OpenStudy (anonymous):

An acute angle θ is in a right triangle with cos θ = 5/6. What is the value of sec θ?

OpenStudy (raden):

hint : sec θ = 1/cos θ

OpenStudy (anonymous):

7

OpenStudy (anonymous):

wow. thats helpful...

OpenStudy (anonymous):

Thanks bro

OpenStudy (raden):

7 not its answer :)

OpenStudy (anonymous):

Yeah it is. Do you even math, bro?

OpenStudy (anonymous):

I don't understand, and yea I know 7 is wrong

OpenStudy (raden):

sec θ = 1/cos θ = 1/(5/6) = 6/5